OpenStudy (anonymous):

Solve the following equation: 9x − (2 − 3x) = 19

OpenStudy (anonymous):

1) distribute the negative sign to each term inside the ( )

OpenStudy (anonymous):

2) Group like terms 3) get x by itself

OpenStudy (anonymous):

\[9x − (2 − 3x) = 19\]\[9x − 2 + 3x = 19\]\[12x − 2= 19\]\[12x = 21\]\[x = {21 \over 12}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

it reduces to 7/4. Do you understand the steps?

OpenStudy (anonymous):

yep it helped alot